What is the even number less thn 85 with the sum of the digits equal to the value of 3 nickels?
Answer by Edwin McCravy(20086)


3 nickels is worth 15 cents. You knew that didn't you?
You know what digits are, don't you?
There are only four two-digit numbers whose digits add up to 15. Here
are three of them.
1. 69 has digits 6 and 9 and 6+9 = 15, and it's less than 85, but it won't do because 69 is odd.
2. 96 has digits 9 and 6 and 9+6 = 15, and it's even, but it won't do because
its greater than 85.
3. 87 has digits 8 and 7 and 8+7 = 15, but that won't do because of two reasons,
not only is it odd but it's also greater than 85
I have left out the answer. See if you can figure it out? There is
only one more integer whose digits add up to 15, and it's even, and it's
also less than 85.
Think real hard and I'm sure you can come up with it.
Edwin